OpenStudy (anonymous):

The density of a substance describes how much matter is present in one unit of the substance's volume. The formula for density is D = m V where m is the mass of the substance and V is the substance's volume. If the density of a substance is equal to 4.1 g/mL and its volume is equal to 9 mL, what is the mass of the substance? A. 13.1 g B. 36.9 g C. 0.46 g D. 2.2 g

OpenStudy (abhisar):

Hello @kaceyk , Welcome to openstudy !

OpenStudy (abhisar):

\(\sf \huge Density = \frac{Mass}{Volume}\)

OpenStudy (abhisar):

So, Masss = Density \(\sf \times\)Volume = 4.1\(\sf \times\)9=36.9 grams

OpenStudy (anonymous):

Thanks and i was trying to fix that

OpenStudy (abhisar):

You probably didn't noticed but you posted your mathematics question in the biology section. Next time check if u r in ryt section.

OpenStudy (abhisar):

ur most welcome :)

OpenStudy (anonymous):

i didnt know that was a math question

OpenStudy (abhisar):

It is :)

OpenStudy (anonymous):

what one